(5x + 2x - 7x – 3) – (6x + 3x – 11)
125x + 4x – 7x – 3 – 36x + 3x – 11
(125x + 4x – 7x – 36x + 3x) – 3 - -11
(125x + 4x + 3x – 43x) – 3 + 11
89x + 8
NOTE: ORIGINAL EQUATION HAD EXPONENTS IN IT:
(5x^3 + 2x^2 - 7x - 3) - (6x^2 + 3x - 11)

Answer by checkley71(8403)

5x^3 does not=125x
(5x)^3=125x^3
(5^3)x=125x
5x^3=5(x^3)
-----------------------------
-----------------------------
5x^3+2x^2-7x-3
-6x^2-3x+11 now add like terms.
-----------------------------------
5x^3-4x^2-10x+7 answer.